Using (x + a) (x + b) = x^2 + (a + b) x + ab, find
103 x 104

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

103 x 104
= (100 + 3)(100 + 4)
= 100^2+ (3 + 4)100 + 12
= 10000 + 700 + 12
= 10712

Thanks